OpenStudy (anonymous):

2x^2+x-1=0 how many real solutions

OpenStudy (ageta):

Let's solve your equation step-by-step. 2x2+x−1=0 Step 1: Factor left side of equation. (2x−1)(x+1)=0 Step 2: Set factors equal to 0. 2x−1=0 or x+1=0

OpenStudy (ageta):

Answer:x=1/2 or x=−1

OpenStudy (anonymous):

if anyone sees the <?> just reload your page...

OpenStudy (ageta):

ok

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

discriminant D=b^2-4ac \[D=1^2-4*2*-1=1+8=9=3^2>0\] roots are real and rational. \[\ because\] it is a qudratic ,so two real roots.